Lenox is planning to read 3 books in 5 weeks. The average number of pages in each of the 3 books is 184. He wants to read the same number of pages everyday except on Friday, when he doesn't plan to read. How many pages should he read per day to finish the 3 books in 5 weeks?
Answer by checkley71(8403)

3*184=552 total pages to read.
5*6=30 days of reading.
552/30=18.4 pages need to be read each day to finish in 5 weeks.